OpenStudy (anonymous):

A poker hand (5 cards) is dealt from a well shuffled deck of cards. Find (a). The chance that the cards are all diamonds; (b). The chance that the cards are all of the same suits.

jimthompson5910 (jim_thompson5910):

There are 13 diamond cards (2, 3, 4, 5, 6, 7, 8, 9, 10, J, Q, K, A) There are 13 C 5 = 1287 ways to deal out 5 diamond cards. There are 52 C 5 = 2598960 ways to deal out 5 cards of any type So the probability of getting 5 diamond cards is 1287/2598960 = 33/66640 = 0.000495

OpenStudy (anonymous):

Thank you so much, I appreciate; but what about the B part of the question?

jimthompson5910 (jim_thompson5910):

When you get the first card, it dictates which suit you're dealing with

jimthompson5910 (jim_thompson5910):

So let's say the first card is a 2 of hearts

jimthompson5910 (jim_thompson5910):

The other four cards must be hearts (to get what you want)

jimthompson5910 (jim_thompson5910):

There are 12 heart cards left There are 4 slots left (in the hand) So there are 12 C 4 = 495 ways to finish up the hand you want There are 51 C 4 = 249900 hands total (after you make your first draw) of any card combo So the probability is 495/249900 = 33/16660 = 0.00198079

jimthompson5910 (jim_thompson5910):

This probability is slightly higher than the one found in part a) So it's more likely to draw 5 cards of the same suit than to just draw all 5 diamond cards. This makes sense because there are more cases in part b), leading to a higher probability.

OpenStudy (anonymous):

Thank you so much.

jimthompson5910 (jim_thompson5910):

np